OpenStudy (lexber):

MEDAL+FAN What is the equation of the line that passes through the points (−2, 1) and (1, 10)? 3x − y = −7 3x − y = −5 3x − y = 5

OpenStudy (anonymous):

slope, m = diffy/diffx m = (1-10)/(-2-1) m = ? y=mx+b

OpenStudy (lexber):

i dont know how to solve it @mthompson440

OpenStudy (anonymous):

What does 1-10= ?

OpenStudy (lexber):

9 @mthompson440

OpenStudy (anonymous):

pay attention to the negative

OpenStudy (lexber):

-9 then @mthompson440

OpenStudy (anonymous):

yes. Now, what does -2-1=?

OpenStudy (lexber):

1!! @mthompson440

OpenStudy (thadds2003):

No, that is wrong, @lexber. It is actually -3, take time and be patient.

OpenStudy (lexber):

sorry @Thadds2003 I am very bad at math... could you help me finish the rest please

OpenStudy (thadds2003):

I kind of don't know this, but lets try from the information that @mthompson gave us. Hold on......

OpenStudy (qwertty123):

Hey I think I know hold on!

OpenStudy (lexber):

okay (:

OpenStudy (thadds2003):

I'm working on the answer......

OpenStudy (lexber):

I understand @Thadds2003 , I like to ave answers from a few people so I learn more!

OpenStudy (thadds2003):

Ok..... Almost done........

OpenStudy (lexber):

Thank you so much @Thadds2003

OpenStudy (thadds2003):

Ok, I might need some help. @Qwertty123 , you said you might know how to do this, correct? If so lets work together.

OpenStudy (qwertty123):

okay

OpenStudy (lexber):

Yes I am willing to work together!! I'm really not that good tbh in math ): bu I'll try hard

OpenStudy (thadds2003):

Okay, so we left of at -3, I need a review on the steps. Anybody willing to get me caught up?

OpenStudy (thadds2003):

WAIT!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (lexber):

I cant bc idk how to do this

OpenStudy (thadds2003):

I HAVE THE ANSWER!!! LET ME TYPE IT OUT!!!

OpenStudy (thadds2003):

I'm 100% sure about this answer......

OpenStudy (thadds2003):

Gradient = (10 - 1) / (1 - (-2)) = 9 / 3 = 3 Using (-2, 1): y = 1 + 3(x - (-2)) = 1 + 3x + 6 = 3x + 7 3x - y = -7 (answer)

OpenStudy (lexber):

really omg thnk you

OpenStudy (lexber):

wow!!! you just saved my life!!! thank yoi so so much

OpenStudy (qwertty123):

Yeah! :)

OpenStudy (thadds2003):

Glad to help. Tag me when you need help, if you like.

OpenStudy (thadds2003):

Are you going to close the question? Medal? lol.

OpenStudy (lexber):

I'm sorry...